# Is Watermarking LLM-Generated Code Robust?

### v2

@@ Abstract
We present the first study of the robustness of existing watermarking techniques on Python code generated by large language models. Although existing works showed that watermarking can be robust for natural language, we show that it is easy to remove these watermarks on code by semantic-preserving transformations.
@@ 1 Introduction
The rapid advancement of large language models (LLMs) like GPT and Codex in understanding and generating code holds transformative potential for software engineering (Chen et al., 2021; OpenAI, 2023; Ugare et al., 2024b). However, it raises concerns about misuse such as code plagiarism and malware generation. Combating misuse requires accurate detection of LLM-generated code, which is challenging as LLMs are designed to produce realistic output that mimics human-generated code.
@@
This Work. We are the first to investigate the robustness of watermarking Python code generated by LLMs. We propose an algorithm that walks the Abstract Syntax Tree (AST) of the watermarked code and randomly applies semantic-preserving program modifications. We observe significantly lower true-positive rate (TPR) of detection even under simple modifications, underscoring the need for robust LLM watermarks tailored specifically for code.
@@ 2 Robustness of Watermarked Code
3 Evaluation
We implement several transformations: InsertDeadCode, Rename, InsertPrint, WrapTryCatch, and Mixed. Appendix A.1 presents details about our experimental setup and the transformations.
@@ 3 Evaluation
We generate Python code completions using the LlamA-7B (Touvron et al., 2023) and CodeLlamA-7B (presented in Appendix A.2) models on the HumanEval (Chen et al., 2021) dataset. We watermark the code with state-of-the-art algorithms UMD and Unigram (Zhao et al., 2023; Kirchenbauer et al., 2023a). We describe these baselines with the hyperparameter values that we used in Appendix A.3. We sequentially apply each program transformation $d=5$ times on the watermarked code. As our transformation procedure is randomized, we run this experiment 3 times and compute the average of the results. Table 1 presents our main results for LlamA-7B. It shows that the program transformations greatly corrupt watermark detectability. Even the simplest transformations InsertPrint and Rename reduce the TPR by at least 1.3x. Complex alterations (e.g., WrapTryCatch and AddDeadCode) reduce the TPR much more significantly.
Refer to captionRefer to caption
Figure 1: TPR vs the number of transformations $d$: UMD (above), Unigram (below).
Varying the Number of Transformations We further show the robustness of the watermark techniques by varying the number of modifications $d$ applied to the watermarked code from 0 to 5. Figure 1 shows that the TPR declines as $d$ increases. For instance, when employing 5 WrapTryCatch modifications, the TPR dropped to 0.22 for the UMD watermark and fell to 0.11 for the Unigram watermark. AddDeadCode and WrapTryCatch modifications exhibit a more pronounced impact on TPR, requiring fewer modifications to reduce TPR by over 2x compared to the other two modifications.
4 Discussion
ACKNOWLEDGMENTS
Appendix A Appendix
A.1 Experimental Setup
Table 2: Watermark detectability results
Transformation Proportion of
tokens changed
AddDeadCode 0.36
InsertPrint 0.12
Rename 0.05
WrapTryCatch 0.27
Mixed 0.22
Table 3: Watermark detectability results
@@ A.1 Experimental Setup
A.2 Evaluation for CodeLLama
Table 3 presents the results of our evaluation on CodeLlamA-7B. Similar to LlamA-7B, we observe that the program transformations greatly reduce watermark detectability. Simple transformations like InsertPrint and Rename reduce the TPR by at least 1.4x. We observe even larger reductions for more complex modifications (e.g., WrapTryCatch and AddDeadCode).
A.3 Watermark Baselines
For both UMD and Unigram, we set $\gamma$ = 0.25, where $\gamma$ represents the fraction of the vocabulary included in the green list. For the Unigram watermark, we set the strength parameter $\delta=2$, where the larger $\delta$ is, the lower the quality of the watermarked LM output, but the easier it is to detect. We observe empirically each function completion only has around 100 tokens on average and the TPR $<0.3$. To increase the number of tokens and thus the TPR, after generating the watermarked code completions, we select 3 random function completions at a time and run detection collectively on the 3 functions. We reject the null hypothesis if $z>3$. We show that the TPR $>0.70$ and the FPR $\leq 0.01$ for the baseline by adopting this approach.
A.4 Related Work
@@
DNN Robustness A large body of research has focused on the robustness of LLMs and other DNNs against adversarial attacks (Zou et al., 2023; Ugare et al., 2024a; Zhang et al., 2023; Laurel et al., 2022; Ugare et al., 2023). This line of work is orthogonal to our investigation as we instead focus on the robustness of whether LLM-generated output (code) can be reliably detected.

### v4

# Is the Watermarking of LLM-Generated Code Robust?

@@ Abstract
We present the first in-depth study on the robustness of existing watermarking techniques applied to code generated by large language models (LLMs). As LLMs increasingly contribute to software development, watermarking has emerged as a potential solution for detecting AI-generated code and mitigating misuse such as plagiarism or the automated generation of malicious programs. While previous research has demonstrated the resilience of watermarking in the text setting, our work reveals that watermarking techniques are significantly more fragile in code-based contexts. Specifically, we show that simple semantic-preserving transformations—such as variable renaming and dead code insertion—can effectively erase watermarks without altering the program’s functionality. To systematically evaluate watermark robustness, we develop an algorithm that traverses the Abstract Syntax Tree (AST) of a watermarked program and applies a sequence of randomized, semantics-preserving transformations. Our experimental results, conducted on Python code generated by different LLMs, indicate that even minor modifications can drastically reduce watermark detectability, with true-positive rates (TPR) dropping below 50% in many cases.
Our code is publicly available at https://github.com/uiuc-arc/llm-code-watermark.
@@ 1 Introduction
The rapid advancement of large language models (LLMs) like GPT and Codex in understanding and generating code holds transformative potential for software development (Chen et al., 2021; OpenAI, 2023; Suresh et al., 2024b; Reddy et al., 2025; Wei et al., 2025b; a; d; c), especially when combined with constrained decoding algorithms (Ugare et al., 2024c; a; Banerjee et al., 2025; Suresh et al., 2025), which improve the syntactic and semantic correctness of the LLM output. However, the use of LLMs for coding tasks raises concerns about misuse such as code plagiarism and malware generation. Combating misuse requires accurate detection of LLM-generated code, which is challenging as LLMs are designed to produce realistic output that mimics human-generated code.
@@
This Work. We are the first to investigate the robustness of watermarking Python code generated by LLMs. We propose an algorithm that walks the Abstract Syntax Tree (AST) of the watermarked code and randomly applies semantic-preserving program modifications. We observe a significantly lower true-positive rate (TPR) of detection even under simple modifications, underscoring the need for robust LLM watermarks tailored specifically for code.
@@ 2.2 Watermarked Program Transformation
3 Experimental Methodology
3.1 Experimental Setup
@@
Transformation Proportion of
tokens changed
AddDeadCode 0.36
InsertPrint 0.12
Rename 0.05
WrapTryCatch 0.27
Mixed 0.22
3.2 Semantic-Preserving Transformations
We implement the following semantic-preserving transformations:
3.3 Watermark Baselines
For both UMD and Unigram, we set $\gamma$ = 0.25, where $\gamma$ represents the fraction of the vocabulary included in the green list. For the Unigram watermark, we set the strength parameter $\delta=2$, where the larger $\delta$ is, the lower the quality of the watermarked LM output, but the easier it is to detect. We observe empirically each function completion only has around 100 tokens on average and the TPR <0.3<0.3. To increase the number of tokens and thus the TPR, after generating the watermarked code completions, we select 3 random function completions at a time and run detection collectively on the 3 functions. We reject the null hypothesis if $z>3$. We show that the TPR >0.70>0.70 and the FPR $\leq 0.01$ for the baseline by adopting this approach.
4 Experimental Results
4.1 Detectability of Watermarked Code For LlamA-7B
Table 2: Llama-7B Watermark detectability results
@@ 4.1 Detectability of Watermarked Code For LlamA-7B
We generate Python code completions using the LlamA-7B (Touvron et al., 2023) on the HumanEval (Chen et al., 2021) dataset. We watermark the code with state-of-the-art algorithms UMD and Unigram (Zhao et al., 2023; Kirchenbauer et al., 2023a). We sequentially apply each program transformation $d=5$ times on the watermarked code. As our transformation procedure is randomized, we run this experiment 3 times and compute the average of the results. Table 2 presents our main results for LlamA-7B. It shows that the program transformations greatly corrupt watermark detectability. Even the simplest transformations InsertPrint and Rename reduce the TPR by at least 1.3x. Complex alterations (e.g., WrapTryCatch and AddDeadCode) reduce the TPR much more significantly.
4.2 Detectability of Watermarked Code For CodeLlamA-7B
Table 3: CodeLlama-7B Watermark detectability results
@@ 4.2 Detectability of Watermarked Code For CodeLlamA-7B
Similarly, we evaluate the detectability of watermarked Python code completions generated by CodeLlamA-7B (Touvron et al., 2023) on HumanEval and present the results in Table 3. Similar to LlamA-7B, we observe that the program transformations greatly reduce watermark detectability. Simple transformations like InsertPrint and Rename reduce the TPR by at least 1.4x. We observe even larger reductions for more complex modifications (e.g., WrapTryCatch and AddDeadCode).
4.3 Varying the Number of Transformations
(a) UMD
(b) Unigram
Figure 1: TPR vs the number of transformations $d$.
We further show the robustness of the watermark techniques by varying the number of modifications $d$ applied to the watermarked code from 0 to 5. Figure 1 shows that the TPR declines as $d$ increases. For instance, when employing 5 WrapTryCatch modifications, the TPR dropped to 0.22 for the UMD watermark and fell to 0.11 for the Unigram watermark. AddDeadCode and WrapTryCatch modifications exhibit a more pronounced impact on TPR, requiring fewer modifications to reduce TPR by over 2x compared to the other two modifications.
5 Related Work
@@
DNN Robustness A large body of research has focused on the robustness of LLMs and other DNNs against adversarial attacks (Suresh et al., 2024a; Tamirisa et al., 2025; Zou et al., 2023; Ugare et al., 2024b; Zhang et al., 2023; Laurel et al., 2022; Ugare et al., 2023). This line of work is orthogonal to our investigation as we instead focus on the robustness of whether LLM-generated output (code) can be reliably detected.
6 Conclusion
ACKNOWLEDGMENTS